# Beyond the Answers: Reviewing the Rationality of Multiple Choice Question Answering for the Evaluation of Large Language Models

### v1

@@ 3.3 Adaptation of MCQA Answer Options
Re-ordering of Answer Options
Previous studies, such as Robinson & Wingate 2022, have demonstrated that re-ordering answer options can induce performance variability, a phenomenon intrinsically linked to the nature of LLMs. For simplicity, we expand each chosen instance into four random distinct orders to assess the sensitivity of LLMs to option sequencing. We conduct experiments with MCQA in its original format on selected instances from MMLU and MedMCQA, recording baseline accuracy (“Acc”) just like the reported accuracy for evaluating LLMs. Subsequently, we introduce a consistency score “$S_{c}$” defined for each instance as follows:
$S_{c}=\frac{\texttt{UNIQUE\_ANSWER}\ (a_{1},\ \dots,\ a_{n})}{n}$
$\bar{S}_{c}=\frac{1}{N}\sum_{i=1}^{N}{S_{c}(i)}$
Variation in Answer Option Quantity
In addition to re-ordered options, our investigation extends to evaluating the influence of varying the number of answer choices. Since the norm of four or five options in most MCQA assessments (Haladyna & Downing, 1993), we deliberate modifications in the number of answer options to examine the adaptability of LLMs to these altered conditions.
When incrementing the number of answer options, we incorporate additional non-semantic tokens as distracting options to circumvent the inadvertent introduction of plausible correct responses, thereby preserving the authenticity of the questions. This expands the number of answer options to 6, 8, and 10 choices. In contrast, when reducing the number of answer options, we randomly eliminate distracting options, simplifying the questions by reducing the options to 3 or 2.
By combining the adaptions of re-ordering and quantity variation of options, Figure 3 demonstrates the performance of LLMs on the MMLU and MedMCQA tasks. The metric “Accuracy” is quantified based on questions with only a single option order for $n$ options. The $\bar{S}_{c}$ and Acc-H represent the mean values quantified across $n$ option orders for $n$ options respectively. Illustratively, taking Baichuan2-13B as an example, it exhibits an accuracy of 70.21% with the original set of four options on MMLU under the zero-shot setting, alongside a notable mean $S_{c}$ of 83.13. However, the Acc-H metric indicates a significant drop, with consistent and correct predictions in merely 39.38% of all instances when each instance is re-ordered into four distinct instances. Observing the performance of Baichuan2-13B with 2 to 10 answer options, a marked volatility in Accuracy is noted, ranging from 44.44% to 75.45%, with the peak accuracy surprisingly achieved with 10 options. Nonetheless, both $S_{c}$ and Acc-H exhibit a declining trend as the number of options increases. This pattern of fluctuating accuracy and diminishing $S_{c}$ and Acc-H values is consistent across all three LLMs on the MMLU. In experiments involving MedMCQA, a similar downward trend is observed for Accuracy, $S_{c}$, and Acc-H with the increment in answer options. Remarkably, for LLaMA2-13B, Acc-H plummets to 0.00% with six or more answer options, casting doubts on the reliability of the LLMs with a larger array of answer options. These observations underscore the questionable reliability of Accuracy metrics when based on a single order of answer options.
For a fine-grained analysis, Table 2 presents the distribution of $S_{c}$ values under the zero-shot setting. It is noticeable that most of the LLMs achieve an $S_{c}$ of at least 3/4. Notably, a higher $S_{c}$ value correlates with higher accuracy in the LLM responses. For instance, the accuracy of ChatGPT on MMLU is 85.23% when $S_{c}=4/4$, but it dwindles to 40.93% at $S_{c}=3/4$ and further plummets to 7.59% at $S_{c}=2/4$. This observed trend, consistent across various settings, suggests that the consistency of predictions across re-ordered instances is indicative of the reliability of the predictions. Consequently, we posit the following hypothesis:
Proposition 3.1.
An LLM can be deemed proficient in answering a specific MCQA-format question only if it consistently predicts the same correct answer across all permutations of option orders.
Table 4: Consistency of LLMs between the MCQA-format and True/False-format questions. correct: True/False questions derived from the correct MCQA options. distracting: True/False questions derived from the distracting MCQA options. † denotes the experiments on the sub-datasets where the LLMs predict correctly with the original MCQA settings.
@@
True or False: ${}^{{\dagger}}$ 0-shot 70.31 70.49 88.20 60.98 87.62 76.58
@@
True or False: ${}^{{\dagger}}$ 0-shot 52.52 55.77 48.51 70.74 68.48 69.87
@@
Altered Option with “None of the Above”
Expanding on previous research conducted with private LLMs (Kadavath et al., 2022), we investigate the impact of introducing a “None of the above” option in place of the correct answer within the test datasets to assess the adaptability of LLaMA, Baichuan, and ChatGPT. As delineated in Table 3, this modification precipitates an extreme decline in performance for the LLMs, with a reduction in both accuracy (Acc) and accuracy-hard (Acc-H) across all models by as much as 70.9%, from 74.44% to a mere 3.54%. Notably, except in the 5-shot scenario for ChatGPT, the models invariably fail to select the “None of the above” option as a substitute for the correct answer, with Acc-H falling below 5% in all other scenarios, underscoring the deficiency in the LLM ability to truly discern the knowledge within the MCQA-format questions.
3.4 Transformation to True-or-False Format
Beyond the alterations in MCQA answer options, we transform the original MCQA-format tasks into a True-or-False format, as depicted in Figure 2. This transformation is exclusively applied to the instances where the LLMs have demonstrated consistent and correct performance (with an $S_{c}$ of $1$). For every MCQA instance with four answer choices, we generate one True-or-False-format question using the correct option and three additional questions using the distracting options, anticipating that the LLMs will respond accurately with “Yes” and “No” respectively. Table 4 elucidates the degree of consistency between the original MCQA-format questions and the adapted True-or-False-format ones. Notably, LLMs exhibit a higher propensity for affirming True-or-False questions that originate from previously selected correct options, evidenced by a consistency rate of up to 91.07%. Conversely, the models demonstrate a reluctance to negate True-or-False questions derived from distracting options. This is evidenced by the fact that up to 51.49% of all distracting options are predicted as correct in the True-or-False format, signifying a divergence in the model decision between MCQA and True-or-False-format questions and further implying the unreliability of MCQA benchmarks.
3.5 Summary
This section has presented a comprehensive experimental analysis, revealing the manifestation of REsponse VAriability Syndrome (REVAS) within Large Language Models. Our empirical findings cast a shadow on the reliability of MCQA as a measure of LLM capabilities, as evidenced by the inconsistent responses of LLMs to derivations of the same MCQA-format question. The identification of REVAS in LLMs is pivotal in recognizing the inherent limitations of current evaluation strategies. Afterward, we will delve into the causative factors for the observed behavior patterns.
4 The More Correct but not the Only Correct.
@@ 3.5 Summary
Figure 4: The propensity of LLMs to select the most correct choice from a set of options and elucidating experiments.
The preceding section has elucidated the enigmatic phenomenon of REVAS, characterized by the unreliable consistency in the responses of LLMs, notably perturbed by minor alterations in the answer options. This inconsistency becomes extreme with the incorporation of the “None of the above” option, often leading to poor performance in MCQA tasks. With these observations, we hypothesize the following:
Assumption 4.1.
As probabilistic generative constructs, LLMs are prone to opt for the most correct option in the MCQA-format task, as opposed to the solely correct option.
Substituting the correct option with “None of the above”, with the above assumption, steers the LLMs towards the selection of the second most correct option, as depicted in Figure 4. To substantiate our assumption, we elaborate experiments encompassing: (1) Option probability analysis; (2) MCQA incorporating multi-selections; and (3) MCQA with the substitution of misleading options, as illustrated in Figure 4. The latter duo of experiments scrutinizes instances where LLMs have erroneously predicted under the original MCQA scenarios.
Table 5: Confidence of specific options in MCQA tasks with LLaMA2-13B and Baichuan2-13B. $C_{c}$: mean confidence of the correct options. $C_{d}$: mean confidence of the most distracting options (with the highest confidence). $C_{d}/C_{c}$: relative confidence score of $C_{d}$ divided by $C_{c}$. † denotes the experiments on the sub-datasets where the LLMs predict correctly with the original MCQA settings.
MMLU MedMCQA
$\textbf{L}^{{\dagger}}$ $C_{c}$ 35.04 35.96
0-shot $C_{d}$ 32.48 30.15
$C_{d}/C_{c}$ 92.69% 83.84%
$C_{c}$ 34.53 35.60
5-shot $C_{d}$ 31.91 30.32
$C_{d}/C_{c}$ 92.41% 85.17%
$\textbf{B}^{{\dagger}}$ $C_{c}$ 65.90 66.32
0-shot $C_{d}$ 61.37 60.08
$C_{d}/C_{c}$ 93.13% 90.59%
$C_{c}$ 65.36 67.30
5-shot $C_{d}$ 61.13 62.12
$C_{d}/C_{c}$ 93.53% 92.30%
@@
In instances where LLMs regard certain distracting options in the MCQA as “correct but not the most correct”, it is plausible for the LLMs to acknowledge these distracting options as correct in other contexts. The aforementioned experiments reveal that question format is a significant factor in model performance. Consequently, we solely consider instances manifesting absolute consistency in all derived instances from the original instance, imposing a more stringent constraint. We probe the confidence of LLMs, as denoted in Chen et al. 2023, by examining the distribution of token probabilities of answer option tokens within the vocabulary list, aiming to discern the existence of “correct” versus “more correct” phenomena. By reformulating MCQA questions into True-or-False questions with both correct and distracting answer options, as represented in Figure 4, we derive independent confidence scores for each option with the probability of token “yes” instead of that of symbol options. Table 5 demonstrates the mean confidence levels of LLMs on the correct ($C_{c}$) and the most distracting ($C_{d}$ the distractor with the highest confidence) options, along with the confidence scores $C_{d}/C_{c}$ to assess the confidence level of $C_{d}$ relatively. Experimental results show that although Baichuan exhibits higher confidence than LLaMA in questions with the correct option, both LLMs attribute a degree of correctness to the most distracting options, with confidence levels exceeding 90% of that associated with the correct options in the MMLU. Consequently, despite accurate predictions and consistency across re-ordered options, LLMs may still perceive certain distracting options as correct, though to a lesser extent compared to the ground-truth options.
Multi-selection in MCQA
Conventionally, LLMs in MCQA paradigms are mostly expected to identify a single correct option. If the most probable option of an LLM diverges from the correct option, erroneous predictions will ensue. Furthermore, as for the wrong cases, if LLMs are prompted to identify all correct options and the correct options are encapsulated within the responded selections, it substantiates the preference of LLMs for “more correct” options over exclusively correct ones. We prompt the LLMs for multiple potential answers, collating instances where the LLMs have made mistakes on the original MCQA tasks. Herein, we designate the distracting options, erroneously identified as correct by the LLMs, as misleading options. Table 6 showcases the recall rates of correct and misleading options in LLM responses under the multi-selection MCQA paradigm, alongside the proportion of instances where LLMs render multiple selections. Results manifest that, in spite of the variance in multi-selection MCQA performance across different LLMs, correct options are encapsulated within the selections in over 78% of instances (up to 92.5%), a metric that surpasses the recall rates for misleading options (only up to 70.31%) previously predicted by the LLMs. This intriguing inconsistency in the recall of correct and misleading options intimates that LLMs not only reclassify unchosen correct options as correct in the multi-selection paradigm but also reassess their stance on misleading options between the single-selection and multi-selection MCQA paradigms.
Table 6: Recall of the correct and misleading options in the responses of LLMs with the setting of multiple selections for MCQA under the few-shot setting. %multi: the proportion of instances where the LLMs generate multiple selections. $R_{correct}$: recall of the correct options in the multiple selections. $R_{misleading}$: recall of the misleading options in the multiple selections.
MMLU MedMCQA
LLaMA2-13B %multi 23.12% 24.38%
$R_{correct}$ 92.50 78.91
$R_{misleading}$ 67.50 70.31
Baichuan2-13B %multi 82.77% 77.68%
$R_{correct}$ 86.29 88.59
$R_{misleading}$ 52.28 48.22
ChatGPT 3.5 %multi 62.14% 68.81%
$R_{correct}$ 85.06 81.27
$R_{misleading}$ 66.67 69.66
MCQA with the Misleading Options Replacement
In line with our definition of misleading options mentioned above, we explore whether replacing these misleading options with arbitrary tokens directs LLMs toward the correct options. This also encompasses instances in which LLMs have misjudged under the original MCQA scenarios. Table 7 elucidates that the LLMs rectify their initial misjudgments in from 39.23% to 57.99% of all instances, evidencing the influence of misleading options on their predictions. These findings also insinuate that certain instances may be intrinsically challenging for LLMs, and substituting a single misleading option may only marginally enhance model performance. The fundamental lack of pertinent knowledge may result in the erroneous predictions of the LLMs.
Table 7: Ratio of the instances where the LLMs turn to predict correctly with the replacement of misleading options under the few-shot setting.
MMLU MedMCQA
LLaMA2-13B 57.99 40.96
Baichuan2-13B 48.55 39.23
ChatGPT 3.5 53.57 44.47
5 LLMs have mostly Learned from the Right
During the stages of pre-training and supervised fine-tuning (SFT), GPT-like LLMs undergo training to predict subsequent text tokens by optimizing the cross-entropy loss function with massive texts. This paradigm predominantly concentrates on the “correct” sequences within textual contexts, given that the bulk of knowledge is mostly encapsulated in assertive statements. This approach, however, may inadvertently neglect a comprehensive understanding of absolute “correctness”, that alternatives to the “correct” answer are inherently incorrect. In light of this observation, we propose the ensuing assumption:
Assumption 5.1.
LLMs predominantly derive knowledge from “correct” contexts encountered during pre-training and supervised fine-tuning, potentially neglecting the learning from “incorrect” contexts.
To empirically evaluate this hypothesis, we employ the Baichuan2-7B-Base model, which is solely pre-trained, and utilize the MedMCQA dataset to curate distinct training and testing sets. Specifically, we extract 10k and 2k non-overlapping instances from MedMCQA, designated as $M_{10k}$ and $M_{2k}$ respectively. The training set comprises: (1) 10k “positive” instances $T_{Pos}$, formulated from questions paired with the correct answers, and (2) 10k “negative” instances $T_{Neg}$, assembled from questions coupled with distracting options. Both subsets, $T_{Pos}$ and $T_{Neg}$, are derived from $M_{10k}$ for the purpose of pre-training. Subsequently, we assemble 2k out-of-distribution instances $T_{SFT}$, extracted from $M_{2k}$, to facilitate the LLMs to learn across diverse question formats during supervised fine-tuning. For the testing phase, we employ an integrated task comprising a balanced amalgamation of original MCQA-format questions, modified MCQA-format questions incorporating the option “None of the above”, and True-or-False-format questions, all sourced from the $M_{10k}$ instances. These testing instances necessitate the same foundational knowledge as those within the $T_{Pos}$ and $T_{Neg}$ sets, albeit presented in varied question formats.
Employing the base model, we execute three distinct training strategies: (1) Only SFT on $T_{SFT}$ to establish a baseline performance. (2) Continuation of pre-training with positive instances $T_{Pos}$, followed by SFT on $T_{SFT}$, to analyse the influence of positive instances. (3) A sequential training involving positive, negative, and SFT instances to ascertain the contribution of negative instances in enhancing model performance. Consistent training steps are maintained across the latter two strategies. The results, as tabulated in Table 8, display that while pre-training only with “positive” instances diminishes performance in both MCQA and True-or-False-format tests, the incorporation of “negative” instances induces a substantial enhancement, particularly in the True-or-False-format test, marked by a 30.53% improvement. This outcome underscores the pivotal role of “negative” instances in fostering a more comprehensive and nuanced learning process within LLMs.
Table 8: Schematic experiments of training the LLMs with positive, negative, and SFT instances on Baichuan2-7B-Base.
Training Strategy MCQA Test True/False Test
SFT 59.81 49.84
Pos + SFT 44.22 (-15.59) 33.01 (-16.83)
Pos + Neg + SFT 68.19 (+8.38) 80.37 (+30.53)
Impact Statements
Appendix A Details of the Performance of the LLMs in Figure 3.

### v2

@@ 3.1 Introducing a “None of the Above” Option
MMLU MedMCQA MMLU MedMCQA MMLU MedMCQA
0-shot Acc 17.17 (-82.83) 12.28 (-87.72) 14.40 (-85.60) 10.32 (-89.68) 23.72 (-76.28) 36.04 (-63.96)
Acc-H 1.01 (-99.99) 0.00 (-100.00) 3.20 (-96.80) 5.56 (-94.44) 8.91 (-91.09) 15.62 (-84.38)
5-shot Acc 6.72 (-93.28) 0.00 (-100.00) 11.72 (-88.28) 4.55 (-95.45) 48.47 (-51.53) 41.67 (-58.33)
Acc-H 0.84 (-99.16) 0.00 (-100.00) 2.34 (-97.66) 0.00 (-100.00) 15.82 (-84.18) 16.67 (-83.33)
3.2 Transforming to True-or-False Format
Next, we transform the original MCQA-format tasks into a True-or-False format. For every MCQA instance with four answer choices, we generate four True-or-False-format questions, including one with the right option (T/F: right) and three questions with the wrong options (T/F: wrong) as depicted in Figure 4, anticipating that the LLMs will respond accurately with “Yes” and “No” respectively. This transformation is exclusively applied to the subsets where the LLMs have demonstrated invariant and correct performance (denoted as $\blacklozenge$).
Figure 4: True-or-false questions derived from the MCQs.
Table 4 provides an analysis of LLM performance on the True-or-False-format questions. Should Question 1 be true, LLMs are expected to exhibit near-perfect accuracy in this context. In practice, while LLMs can achieve an accuracy of up to 91.07% on the T/F: right datasets, the models demonstrate a reluctance to negate statements in the True-or-False questions derived from wrong options. This is evidenced by the fact that up to 51.49% of T/F: wrong datasets are predicted as right in the True-or-False format, signifying a divergence in the model decision between MCQA and True-or-False-format questions and implying the unreliability of MCQA benchmarks.
Table 4: Performance of LLMs on the True-or-False questions. $\blacklozenge$ means the subsets of datasets where LLMs have answered correctly across all re-ordered answer options of the MCQs.
@@ 3.2 Transforming to True-or-False Format
T/F: right 0-shot 70.31 70.49 88.20 60.98 87.62 76.58
@@
T/F: wrong 0-shot 52.52 55.77 48.51 70.74 68.48 69.87
@@
The experiments conducted have yielded the following insights:
Observation 3: While LLMs can display invariability on specific MCQs, their performance noticeably declines when confronted with modified MCQs under the following scenarios: (1) the correct answer options are replaced by “None of the above” and (2) dealing with True-or-False questions formulated from incorrect answer options.
3.3 The Most Correct One but not the Only Correct One.
Observation 3 suggests that LLMs may generate unexpected predictions when presented with modified questions lacking the original right answer option from the corresponding MCQs. Consequently, we can formulate the following intuitive Assumption 1, as illustrated in Figure 5. To validate Assumption 1, we elaborate a series of experiments.
Assumption 1: While LLMs display invariability on specific MCQs with a consistent answer option, they may regard this option as the most accurate, but not exclusively so. Specifically, LLMs might acknowledge the potential correctness of other wrong options, though not more correct than the chosen options.
@@
Figure 5: Illustration of Assumption 1: LLMs may just select the most correct choice from an option set.
Figure 6: Experiments validating Assumption 1. “Option probability analysis” investigates the MCQs that LLMs consistently and accurately predict on all re-ordered orders. “Multiple selections” and “Replacement of misleading options” focus on the MCQs that LLMs erroneously predict.
@@
Utilizing MMLU and MedMCQA, we investigate the confidence of LLMs, as outlined in [chen-etal-2023-close], by analyzing the distribution of token probabilities for answer option tokens. By converting MCQs into True-or-False questions with both right and wrong answer options, as represented in Figure 6, we derive the confidence scores based on the token probability for each answer option. Table 5 demonstrates the mean confidence of LLMs for the right options ($\text{Conf}_{right}$) and the wrong options ($\text{Conf}_{wrong^{\ast}}$, the wrong option that scores with the highest confidence in all wrong options for each MCQ), along with the relative confidence scores. The experimental results show that while LLMs consistently consider the wrong options as less correct than the right options, demonstrated by all relative confidence being below 100%, the wrong options still achieve substantial confidence ranging from 83.84% to 92.69% of those for the right options. Consequently, despite accurate predictions and invariability across re-ordered options, LLMs may still perceive certain wrong options as correct, though to a lesser extent compared to the right ones.
Table 5: Confidence of answer options in MCQA tasks with open-source LLMs. $\text{Conf}_{right}$ : mean confidence of the right options. $\text{Conf}_{wrong^{\ast}}$: mean confidence of the wrong options that score with the highest confidence. $\text{Conf}_{wrong^{\ast}}/\text{Conf}_{right}$: relative confidence score. $\blacklozenge$ denotes the experiments on the sub-datasets where the LLMs predict correctly with the original MCQA settings.
LLaMA2-13B Baichuan2-13B
MMLU MedMCQA MMLU MedMCQA
$\text{Conf}_{right}$ 35.04 35.96 65.90 66.32
0-shot $\text{Conf}_{wrong^{\ast}}$ 32.48 30.15 61.37 60.08
$\text{Conf}_{wrong^{\ast}}/\text{Conf}_{right}$ 92.69% 83.84% 93.13% 90.59%
$\text{Conf}_{right}$ 34.53 35.60 65.36 67.30
5-shot $\text{Conf}_{wrong^{\ast}}$ 31.91 30.32 61.13 62.12
$\text{Conf}_{wrong^{\ast}}/\text{Conf}_{right}$ 92.41% 85.17% 93.53% 92.30%
MCQA with Multiple Selections
For the MCQA datasets involved in this study, LLMs are tasked with identifying only one correct option per MCQ. We collect the instances where LLMs incorrectly predict the answers, denoted as MMLU and MedMCQA. In the above instances, the wrong options which LLMs have regarded as the right ones mistakenly are defined as misleading options. Then, LLMs are prompted to recognize all plausible correct options among all answer options. Table 6 showcases the recall for the right and misleading options, along with the proportion of instances where LLMs render multiple selections. The results reveal that the right options are included in the selections in over 78% of instances, reaching up to 92.5%. This indicates that the LLMs also recognize the right options as correct but less correct than the misleading ones.
[…]
LLaMA2-13B Baichuan2-13B ChatGPT 3.5
MMLU MedMCQA MMLU MedMCQA MMLU MedMCQA
Multi-selections %multi 23.12% 24.38% 82.77% 77.68% 62.14% 68.81%
Recallr 92.50 78.91 86.29 88.59 85.06 81.27
Recallm 67.50 70.31 52.28 48.22 66.67 69.66
Replacement Accuracy 57.99 40.96 48.55 39.23 53.57 44.47
MCQA with the Misleading Option Replacement
Apart from the multi-selection scenario, we explore the impact of replacing misleading options with arbitrary non-semantic tokens in MMLU and MedMCQA. Table 6 elucidates that the LLMs correctly identify the right options in 39.23% to 57.99% of instances, highlighting the influence of misleading options on their predictions. For cases where LLMs continue to make incorrect predictions, a fundamental deficit in relevant knowledge likely underpins the LLM incapability to generate the correct answers.
The analyses conducted across the three experimental scenarios support the likelihood of Assumption 1 being valid. This leads to an observation that underscores a fundamental limitation in using MCQA-based evaluations to accurately gauge the capabilities of LLMs:
Observation 4: In the context of MCQA, while LLMs may select the correct answer, there is a possibility that they also recognize other wrong options as correct.
4 MCQA+: Enhancing MCQA for Robust LLM Evaluation
Experimental analyses have revealed the limitations of the MCQA benchmark for evaluating LLMs. In response, we propose an augmenting approach based on the original MCQA dataset referring to empirical findings in the above observations, termed MCQA+. The MCQA+ dataset comprises (1) the original MCQs; (2) MCQs with re-ordered answer options; (3) MCQs featuring a different number of answer options; (4) MCQs where the correct options are replaced with “None of the above”; and (5) True-or-False questions derived from both right and wrong answer options. We convert each instance from the original MCQA format to the MCQA+ format and adopt the accuracy as the evaluation metric for LLMs (details provided in Appendix B).
Table 7 illustrates the comparative performance on the original MCQA dataset and the MCQA+ dataset. For LLaMA2 and Baichuan2, performance on the MMLU benchmark shows a substantial decline in both the 0-shot and 5-shot scenarios. For instance, accuracy for LLaMA2 plummeted from 72.73% to 45.21% in the 0-shot setting. This suggests that LLaMA2 and Baichuan2 might have been overfitted to the widely-used MMLU benchmark. For ChatGPT, a performance decrease is also evident under the 5-shot setting on MMLU. In contrast, performance on the MedMCQA dataset remains relatively stable across all three LLMs, possibly due to its specialized medical context. Notably, while LLaMA2 and Baichuan2 initially outperform ChatGPT on the original MMLU dataset, ChatGPT achieves markedly higher accuracy with MCQA+, suggesting its utility in revealing true model capabilities.
Although the MCQA+ evaluation provides a more reliable measure of LLM capabilities, it entails considerably higher computational costs compared to the original MCQA. To address this, we implement a sampling method for each MCQ instance, randomly selecting one from the expanded MCQA+ instances to form MCQA+ (×1)($\times$ 1)( × 1 ), which requires the same computational cost as the original MCQA. Table 7 confirms that even with this cost-efficient approach, MCQA+ (×1)($\times$ 1)( × 1 ) effectively exposes true capabilities.
Table 7: Model performance on the original MCQA dataset and MCQA+ with the accuracy metric.
LLaMA2-13B Baichuan2-13B ChatGPT 3.5
MMLU MedMCQA MMLU MedMCQA MMLU MedMCQA
0-shot MCQA 72.73 43.23 70.21 45.81 65.00 56.88
MCQA+ 45.21 (-27.52) 41.20 (-2.02) 55.53 (-14.68) 47.52 (+1.71) 67.02 (+2.02) 56.50 (-0.38)
MCQA+ (×$\times 1$) 46.63 (-26.10) 40.07 (-3.16) 55.67 (-14.55) 44.71 (-1.10) 66.13 (+1.13) 56.00 (-0.88)
5-shot MCQA 74.44 43.33 71.26 43.95 89.80 59.77
MCQA+ 57.46 (-16.98) 42.37 (-0.96) 54.65 (-16.61) 45.13 (+1.18) 70.24 (-19.56) 61.04 (+1.27)
MCQA+ (×$\times 1$) 57.85 (-16.59) 40.88 (-2.45) 54.13 (-17.13) 45.89 (+1.94) 71.44 (-18.36) 61.03 (+1.26)
5 Related Work
Transformer-based language models [vaswani2017attention] have led the research of NLP into a new era. Prior work has elucidated the scaling law, positing that the efficacy of language models is typically proportional to both the model size and the magnitude of pre-training data [henighan2020scaling, kaplan2020scaling]. This principle has catalyzed a paradigm shift towards LLMs, with the number of parameters escalating from millions [bert, gpt2] to billions [radford2018improving, gpt3]. Recent advancements, including supervised fine-tuning and alignment with human values [instruct, chung2022scaling, bai2022training], have further augmented the capabilities of LLMs, enabling them to adhere more closely to human instructions and ethical considerations. Nonetheless, challenges persist since LLMs may show variability in the model responses, especially under the scenarios of MCQA. [robinson2022leveraging] termed the ability to associate the answer options and corresponding symbols as multiple choice symbol binding (MCSB) and proved that the MCSB ability varied significantly by models. Additionally, [wang2023large] revealed vulnerabilities in the ranking of candidate responses, which could be manipulated by altering the presentation order. [zheng2023large] investigated the token bias in LLMs, which could lead to selection biases during option prediction. Furthermore, [kadavath2022language] explored the reliability of the LLM performance and calibration, focusing exclusively on a set of private models under the MCQA settings. Their methodology involved modifying the options and transforming the original questions into binary-choice questions answering with options of “True” and “False”. Building on these foundational studies, in this study we conduct a comprehensive range of experiments applied to a diverse series of LLMs, including both open-source and commercial models. We aim to assess their performance invariability across various settings while using the same set of instances, thereby providing a broader understanding of their capabilities and limitations in different application scenarios.
6 Conclusion
In this study, we critically examined the efficacy of MCQA as a metric for evaluating LLMs, identifying key limitations of MCQA in capturing the true model capacities of LLMs. Our analysis revealed a tendency for overfitting in the context of four-option MCQs and highlighted potential behavioral logic issues that LLMs encounter with MCQs. The findings suggest that MCQA, in its conventional form, does not adequately reflect the full capacities of LLMs, as models may incorrectly recognize wrong options as correct. To overcome these deficiencies, we introduced MCQA+, an enhanced version of the original MCQA, designed to more accurately reveal the true capabilities of LLMs and improve the reliability of evaluation metrics.
[…]
Appendix A Variability of LLMs
Ideally, an intelligent LLM should display invariability to the re-ordered answer options for a specific MCQ, since the knowledge required by the question remains unchanged. We conduct experiments with MCQA in its original format on selected instances from MMLU and MedMCQA, recording baseline accuracy (“Acc”) just like the reported accuracy for evaluating LLMs. For a MCQ $q$, it includes four answer options $(opt_{1},opt_{2},opt_{3},opt_{4})$. We randomly permute the answer options into four orders, like:
$q_{1}=(q,A:opt_{1},B:opt_{2},C:opt_{3},D:opt_{4})$
$q_{2}=(q,A:opt_{2},B:opt_{4},C:opt_{1},D:opt_{3})$
$q_{3}=(q,A:opt_{2},B:opt_{1},C:opt_{3},D:opt_{4})$
$q_{4}=(q,A:opt_{4},B:opt_{1},C:opt_{3},D:opt_{2})$
LLMs are tested with the above questions with re-ordered answer options and generate answer options. Let $S=(opt_{i},opt_{j},\dots)$ be the set of answer options.
For $opt_{k}$ in $S$, $f(opt_{k})$ is the frequency of $opt_{k}$ in $S$. Subsequently, we introduce a consistency score “$S_{c}$” defined as follows:
$S_{c}=\frac{\max_{opt_{k}\in S}f(opt_{k})}{n}$ (1)
$\bar{S}_{c}=\frac{1}{N}\sum_{i=1}^{N}{S_{c}(i)}$ (2)
Apart from the analysis in Section 2.2.1, by combining the $S_{c}$, Figure 7 demonstrates a deeper analysis. The metric “Accuracy” is quantified based on questions with only a single option order for $n$ options. The $\bar{S}_{c}$ and Acc-H represent the mean values quantified across $n$ option orders for $n$ options respectively. Illustratively, taking Baichuan2-13B as an example, it exhibits an accuracy of 70.21% with the original set of four options on MMLU under the zero-shot setting, alongside a notable mean $S_{c}$ of 83.13. However, the Acc-H metric indicates a significant drop, with consistent and correct predictions in merely 39.38% of all instances when each instance is re-ordered into four distinct instances. Observing the performance of Baichuan2-13B with 2 to 10 answer options, a marked volatility in Accuracy is noted, ranging from 44.44% to 75.45%, with the peak accuracy surprisingly achieved with 10 options. Nonetheless, both $S_{c}$ and Acc-H exhibit a declining trend as the number of options increases. This pattern of fluctuating accuracy and diminishing $S_{c}$ and Acc-H values is consistent across all three LLMs on the MMLU. In experiments involving MedMCQA, a similar downward trend is observed for Accuracy, $S_{c}$, and Acc-H with the increment in answer options. Remarkably, for LLaMA2-13B, Acc-H plummets to 0.00% with six or more answer options, casting doubts on the reliability of the LLMs with a larger array of answer options. These observations underscore the questionable reliability of Accuracy metrics when based on a single order of answer options. Table 8 displays detailed results in Figure 7.